# Specific heat study

### v1

@@ Abstract
We examine the $S=1/2$ Heisenberg magnet on four three-dimensional lattices – simple-cubic, diamond, pyrochlore, and hyperkagome ones – for ferromagnetic and antiferromagnetic signs of the exchange interaction in order to illustrate the effect of lattice geometry on the finite-temperature thermodynamic properties with a focus on the specific heat $c(T)$. To this end, we use quantum Monte Carlo simulations or high-temperature expansion series complemented with the entropy method. We also discuss a recent proposal about hidden energy scale in geometrically frustrated magnets.
quantum Heisenberg spin model, geometrically frustrated lattices
pacs:
75.10.Jm
Its our honor and pleasure to contribute to the Condensed Matter Physics journal project celebrating I.R. Yukhnovskii’s 100th birthday. Two of us (T.K. and O.D.) were strongly influenced by his personality since 1980s. We keep in mind so many lovely memories of him, and not only those related to physics but also to life in general. Looking back over those years, we realize how green and simple-minded we were and how good were his advises. We acknowledge the kindness, that always came out from him, with deep gratitude.
I Introduction
Statistical mechanics and thermodynamics are of great importance for understanding the basic nature of the world. Thus, the quantum era was opened by solving three problems, two of which – the black-body radiation (M. Planck) and the specific heat of solids (A. Einstein) – concern statistical mechanics and thermodynamics [1]. As a matter of fact, understanding the specific heat, i.e., the amount of heat required to change the temperature of an electromagnetic field or a material by unit, broke the boundaries of traditional physics.
@@ I Introduction
$H=J\sum_{\langle mn\rangle}\bm{S}_{m}\cdot\bm{S}_{n}.$ (1)
Here, $J=1$ corresponds to the antiferromagnetic exchange interaction whereas J=1J=-1 corresponds to the ferromagnetic exchange interaction and $\langle mn\rangle$ denotes the neighboring lattice sites $m$ and $n$. Furthermore, spin-1/2 operators obey the following commutation relations:
$\left[S_{m}^{x},S_{n}^{y}\right]={\rm i}\delta_{mn}S_{m}^{z}$ (2)
($\hbar=1$) and two more similar relations with the cyclically shifted order of $x,y,z$. Obviously, Hamiltonian (1) is a sum of noncommuting terms and therefore represents an essentially quantum system.
@@
$Z(T,N)=\sum_{j=1}^{2^{N}}\exp\left(-\frac{E_{j}}{T}\right)$ (3)
@@
$f(T)=\lim_{N\to\infty}\frac{F(T,N)}{N},\;F(T,N)=-T\ln Z(T,N).$ (4)
In particular, for the entropy $s$ and the specific heat $c$ we have:
$s(T)=-\frac{\partial f}{\partial T},\;\;\;c(T)=T\frac{\partial s(T)}{\partial T}.$ (5)
@@
Since the set of eigenvalues $\{E_{j}/J\}$ is identical for ferromagnetic and antiferromagnetic cases, e.g., for $N=2$ we have a singlet with E/J=3/4E/J=-3/4 and a triplet with $E/J=1/4$, the partition function (3) for $J=1$ and $T>0$ equals to the partition function (3) for J=1J=-1 and $T<0$. However, it is not clear how this relation can be used in practice. In this study, we consider both signs of the exchange coupling, $J>0$ and $J<0$.
@@
It may be instructive to bear in mind a simpler case of the Ising interaction in Eq. (1) when $\bm{S}_{m}\cdot\bm{S}_{n}\to S_{m}^{z}S_{n}^{z}$. Then all terms in the Hamiltonian commute and, from such a perspective, one faces a classical system. Consider the Ising model on a bipartite lattice (like square, simple-cubic etc.), which consists of two sublattices and all sites of one sublattice are surrounded by the sites of the other sublattice and vice versa. A rotation of the axes in the spin space for one sublattice results in the change of the Ising interaction sign. Therefore, the thermodynamics (but not correlations) does not depend on the sign of $J$. For an order-disorder phase transition this means that the Curie temperature equals the Néel temperature, i.e., $T_{\rm C}=T_{\rm N}$ as for the square-lattice Ising model $T_{\rm C}=T_{\rm N}\approx 0.567$ [2, 3] or for the honeycomb-lattice Ising model $T_{\rm C}=T_{\rm N}\approx 0.380$ [3]. Frustrated Ising lattices are different. For example, the triangular-lattice ferromagnetic Ising model has the Curie temperature $T_{\rm C}\approx 0.910$ [3], whereas its antiferromagnetic counterpart shows no order even at $T=0$ but, instead, has an exponentially large ground-state entropy [4, 5].
@@
Recently, some interesting relations between the Heisenberg and Ising models on a frustrated lattice have been suggested in Refs. [6, 7]. The authors of these papers argued that the specific heat $c(T)$ of the frustrated-lattice Heisenberg model like kagome, hyperkagome etc. illustrates a hidden energy scale (which is significantly lower than $J$) as an extra low-temperature peak. Moreover, the accumulated entropy equals to the residual ground-state entropy of the Ising model on the same lattice.
@@
In this paper, we investigate the $S=1/2$ Heisenberg model on four three-dimensional lattices, see Fig. 1. Two of them, simple-cubic and diamond, are bipartite ones, whereas other two, pyrochlore and hyperkagome, are geometrically frustrated ones. All these quantum spin-lattice systems are widely used in the theory of magnetism. Our goal is to discuss thermodynamic quantities for these models focusing on the specific heat in order to highlight the effect of lattice geometry in the features of $c(T)$ for the quantum Heisenberg ferro- and antiferromagnets. To this end, we use quantum Monte Carlo simulations (ferromagnets and bipartite-lattice antiferromagnets) and high-temperature expansion series analysis (frustrated-lattice antiferromagnets). We have to notice that some of the presented results are not completely new but are scattered over literature, and here we have collected them in one place accompanied by not reported yet results and they all are considered from a common perspective.
@@
The rest of the paper is organized as follows. In Sections II, we introduce the models and methods. Then, in Section III, we report our findings focusing on the temperature profiles of the specific heat. Finally, we discuss and summarize in Section IV. Some additional information on the entropy-method calculations for the pyrochlore and hyperkagome cases is presented in Appendix A and Appendix B, respectively.
II Set up the problem
II.1 Lattices
@@
Figure 1: Simple-cubic, diamond, pyrochlore, and hyperkagome lattices which are considered in this paper. For a detailed description see the main text, Sec. II.1.
@@
Let us describe the lattices to be considered, see Fig. 1. Such descriptions are required for quantum Monte Carlo simulations using the ALPS package [8, 9] as well as for generating high-temperature expansion series [10].
@@
The sites on the simple-cubic lattice are defined by $\bm{R}_{\bm{m}}=n_{x}\bm{e}_{x}+n_{y}\bm{e}_{y}+n_{z}\bm{e}_{z}$, where $\bm{e}_{x}=(1,0,0)$, $\bm{e}_{y}=(0,1,0)$, $\bm{e}_{z}=(0,0,1)$ and $n_{x}$, $n_{y}$, $n_{z}$ are integers. Here and below we set the length of the cube edge to unity.
@@
The diamond lattice consists of two interpenetrating face-centered-cubic sublattices displaced by a vector $(1,1,1)/4$. That is, the sites on the diamond lattice are defined by $\bm{R}_{\bm{m}\alpha}=\bm{R}_{\bm{m}}+\bm{r}_{\alpha}$. The sites of the face-centered-cubic lattice are determined by $\bm{R}_{\bm{m}}=m_{1}{\bm{e}}_{1}+m_{2}{\bm{e}}_{2}+m_{3}{\bm{e}}_{3}$, where ${\bm{e}}_{1}=(0,1,1)/2$, ${\bm{e}}_{2}=(1,0,1)/2$, ${\bm{e}}_{3}=(1,1,0)/2$ and $m_{1}$, $m_{2}$, $m_{3}$ are integers; the origins of the two face-centered-cubic sublattices are taken to be $\bm{r}_{1}=(0,0,0)$ and $\bm{r}_{2}=(1,1,1)/4$.
@@
The pyrochlore lattice consists of four interpenetrating face-centered-cubic sublattices with the lattice sites defined by $\bm{R}_{\bm{m}\alpha}=\bm{R}_{\bm{m}}+\bm{r}_{\alpha}$. Here the sites of the face-centered-cubic lattice are defined above, whereas the origins of the four face-centered-cubic sublattices are taken to be $\bm{r}_{1}=(0,0,0)$, $\bm{r}_{2}=(0,1,1)/4$, $\bm{r}_{3}=(1,0,1)/4$, and $\bm{r}_{4}=(1,1,0)/4$.
@@
Finally, the hyperkagome lattice has 12 sites in a cubic unit cell. The sites on the hyperkagome lattice are defined by $\bm{R}_{\bm{m}\alpha}=\bm{R}_{\bm{m}}+\bm{r}_{\alpha}$. Here, $\bm{R}_{\bm{m}}=n_{x}\bm{e}_{x}+n_{y}\bm{e}_{y}+n_{z}\bm{e}_{z}$ generates a simple-cubic-lattice translation (see above) and the origins of twelve equivalent sublattices are defined by $\bm{r}_{\alpha}$, $\alpha=1,2,\dots,12$, where $\bm{r}_{1}=(-2,0,2)/4$, $\bm{r}_{2}=(-1,3,2)/4$, $\bm{r}_{3}=(-2,3,1)/4$, $\bm{r}_{4}=(-1,1,0)/4$, $\bm{r}_{5}=(-2,1,3)/4$, $\bm{r}_{6}=(-1,2,3)/4$, $\bm{r}_{7}=(-3,2,1)/4$, $\bm{r}_{8}=(0,2,2)/4$, $\bm{r}_{9}=(0,1,1)/4$, $\bm{r}_{10}=(-3,3,0)/4$, $\bm{r}_{11}=(0,0,0)/4$, and $\bm{r}_{12}=(-3,0,3)/4$.
@@
II.2 Methods
II.2.1 Quantum Monte Carlo
@@
II.2.2 High-temperature expansion series augmented by the entropy method
Using Ref. [10], we have generated the high-temperature expansion series for the specific heat (per site) $c(\beta)$, $\beta=1/T$,
$c(\beta)=\sum_{i=2}^{n}d_{j}\beta^{i}.$ (6)
Here, for the pyrochlore lattice $n=17$, the coefficients $d_{2},\ldots,d_{13}$ were reported in Ref. [13] and $d_{14},\ldots,d_{17}$ in Ref. [14]; for the hyperkagome lattice $n=20$ and the coefficients $d_{17},\ldots,d_{20}$,
@@ II.2.2 High-temperature expansion series augmented by the entropy method
complement those that were reported in Ref. [15], see also Appendix B.
@@
Within the entropy method [16, 17, 18], one works with the entropy as a function of the energy $s(e)$, $e_{0}\leq e\leq 0$, $e_{0}$ is the ground-state energy, and the Maclaurin series for $s(e)$ is known from Eq. (6). Assuming gapples low-energy excitations (they yield a power-law decay of the specific heat $c(T)$ as $T\to 0$ with the exponent $\alpha$), one interpolates an auxiliary function $G(e)=[s(e)]^{(1+\alpha)/\alpha}/(e-e_{0})$ between $e=e_{0}$ and $e=0$ using Padé approximants $[u,d](e)$ with $d+u\leq n$, $n=17$ (pyrochlore) or $n=20$ (hyperkagome): $G(e)\approx[(\ln 2)^{\alpha/(1+\alpha)}/(-e_{0})]\,[u,d](e)$. As a result one gets the approximate entropy $s(e)\approx[(e-e_{0})G(e)]^{\alpha/(1+\alpha)}$ and therefore, the specific heat $c(T)$ in the parametric form: $T=1/s^{\prime}(e)$, $c(e)=-[s^{\prime}(e)]^{2}/s^{\prime\prime}(e)$; here $(\ldots)^{\prime}={\rm d}(\ldots)/{\rm d}e$. Thus, besides high-temperature expansion series (6), one needs the values of $e_{0}$ and $\alpha$. Under another assumption, when low-energy excitations are gapped, one works with another auxiliary function $G(e)=(e-e_{0})[s(e)/(e-e_{0})]^{\prime}\approx(\ln 2/e_{0})[u,d](e)$, which yields the approximate entropy $s(e)/(e-e_{0})\approx-\ln 2/e_{0}-\int_{e}^{0}{\rm d}\xi G(\xi)/(\xi-e_{0})$, and one needs as an input the value of $e_{0}$ only.
@@
It should be noted that even if the required input information is unavailable, one may still proceed with the entropy method and find those parameters that are missing searching for such their values which yield the largest number of close to each other $c(T)$ based on various Padé approximants $[u,d](e)$ [19]. Such a heuristic criteria was used in 2020 [13] in the study of the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet on the basis of 13th order high-temperature expansion series to arrive at $e_{0}\approx-0.52$ and $\alpha=2$ as a most favorable set of the unknown parameters. For the hyperkagome-lattice case, similar arguments lead to $e_{0}$ which lies between 0.441-0.441 and 0.435-0.435 and $\alpha=2$ [20].
@@
Since 2021, there has been several proposals for the ground-state energy of the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet: 0.490(6)-0.490(6) (density-matrix renormalization group method [21]), 0.477(3)-0.477(3) (variational Monte Carlo methods [22]). Besides, more terms in high-temperature expansion series have become available now [14]. Furthermore, for the hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnet, four more coefficients in Eq. (6) are available now in comparison to Ref. [15], see Eq. (II.2.2). All this motivates us to reconsider the entropy-method calculations for the pyrochlore- and hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnets and refine the temperature profile $c(T)$, see Sec. III.2.2 below.
III Specific heat $c(T)$

### v3

# Three-dimensional unfrustrated and frustrated quantum Heisenberg magnets. Specific heat study

@@ Abstract
We examine the $S=1/2$ Heisenberg magnet on four three-dimensional lattices — simple-cubic, diamond, pyrochlore, and hyperkagome ones — for ferromagnetic and antiferromagnetic signs of the exchange interaction in order to illustrate the effect of lattice geometry on the finite-temperature thermodynamic properties with a focus on the specific heat $c(T)$. To this end, we use quantum Monte Carlo simulations or high-temperature expansion series complemented with the entropy method. We also discuss a recent proposal about hidden energy scale in geometrically frustrated magnets.
Key words: quantum Heisenberg spin model, geometrically frustrated lattices
Abstract
Ми дослiджуємо $S=1/2$ магнетик Гайзенберга на чотирьох тривимiрних гратках — простiй кубiчнiй, алмазу, пiрохлору i гiперкагоме — для феромагнiтного i антиферомагнiтного знакiв обмiнної взаємодiї, щоб проiлюструвати ефект геометрiї гратки на скiнченнотемпературнi термодинамiчнi властивостi, зокрема, на теплоємнiсть $c(T)$. З цiєю метою ми використовуємо симуляцiї методом квантового Монте Карло i високотемпературнi розвинення, доповненнi методом ентропiї. Ми також обговорюємо недавню пропозицiю про прихований енергетичний масштаб у геометрично фрустрованих магнетиках.
Ключовi слова: квантова Гайзенбергова спiнова модель, геометрично фрустрованi гратки
Its our honor and pleasure to contribute to the Condensed Matter Physics journal project celebrating I. R. Yukhnovskii’s 100th birthday. Two of us (T. K. and O. D.) were strongly influenced by his personality since 1980s. We keep in mind so many lovely memories of him, and not only those related to physics but also to everyday life in general. Looking back over those years, we realize how green and simple-minded we were and how good were his pieces of advise. With deep gratitude we acknowledge the kindness that always irradiated from him.
1 Introduction
Statistical mechanics and thermodynamics are of great importance for understanding the basic nature of the world. Thus, the quantum era was opened by solving four problems, two of which — the black-body radiation (M. Planck) and the specific heat of solids (A. Einstein) — concern statistical mechanics and thermodynamics [1, 2, 3]. As a matter of fact, understanding a specific heat, i.e., the amount of heat required to change the temperature of an electromagnetic field or a material in equilibrium with a bath by unit, broke the boundaries of traditional physics.
@@ 1 Introduction
$H=J\sum_{\langle mn\rangle}{\bf{S}}_{m}\cdot{\bf{S}}_{n}\,.$ (1.1)
Here, $J=1$ corresponds to the antiferromagnetic exchange interaction whereas J=1J=-1 corresponds to the ferromagnetic exchange interaction and $\langle mn\rangle$ denotes the neighboring lattice sites $m$ and $n$. Furthermore, spin-1/2 operators ${\bf{S}}_{m}=(S_{m}^{x},S_{m}^{y},S_{m}^{z})$ obey the following commutation relations:
$\left[S_{m}^{x},S_{n}^{y}\right]={\rm i}\delta_{mn}S_{m}^{z}$ (1.2)
($\hbar=1$) and two more similar relations with the cyclically shifted order of $x$, $y$, $z$. Obviously, Hamiltonian (1.1) is a sum of noncommuting terms and therefore represents an essentially quantum system.
@@
$Z(T,N)=\sum_{j=1}^{2^{N}}\exp\left(-\frac{E_{j}}{T}\right)$ (1.3)
@@
$f(T)=\lim_{N\to\infty}\frac{F(T,N)}{N},\quad F(T,N)=-T\ln Z(T,N).$ (1.4)
In particular, for the entropy per site $s$ and the specific heat per site $c$ we have:
$s(T)=-\frac{\partial f}{\partial T},\quad c(T)=T\frac{\partial s(T)}{\partial T}.$ (1.5)
@@
Since the set of eigenvalues $\{E_{j}/J\}$ is identical for ferromagnetic and antiferromagnetic cases, e.g., for $N=2$ we have a singlet with E/J=3/4E/J=-3/4 and a triplet with $E/J=1/4$, the partition function (1.3) for $J=1$ and $T>0$ equals to the partition function (1.3) for J=1J=-1 and $T<0$. However, it is not clear how this relation can be used in practice. In this study, we consider both signs of the exchange coupling, $J>0$ and $J<0$.
@@
It may be instructive to bear in mind a simpler case of the Ising interaction in equation (1.1) when ${\bf{S}}_{m}\cdot{\bf{S}}_{n}\to S_{m}^{z}S_{n}^{z}$. Then, all terms in the Hamiltonian commute and, from such a perspective, one faces a classical system. Consider the Ising model on a bipartite lattice (such as square, simple-cubic etc.), which consists of two sublattices and all sites of one sublattice are surrounded by the sites of the other sublattice and vice versa. A rotation of the axes in the spin space for one sublattice results in the change of the Ising interaction sign. Therefore, the thermodynamics (rather than correlations) does not depend on the sign of $J$. For an order-disorder phase transition this means that the Curie temperature equals the Néel temperature, i.e., $T_{\rm C}=T_{\rm N}$ as for the square-lattice Ising model $T_{\rm C}=T_{\rm N}\approx 0.567$ [4, 5] or for the honeycomb-lattice Ising model $T_{\rm C}=T_{\rm N}\approx 0.380$ [5]. Frustrated Ising lattices are different. For example, the triangular-lattice ferromagnetic Ising model has the Curie temperature $T_{\rm C}\approx 0.910$ [5], whereas its antiferromagnetic counterpart shows no order even at $T=0$ but, instead, has an exponentially large ground-state entropy [6, 7].
@@
Recently, some interesting relations between the Heisenberg and Ising models on a frustrated lattice have been suggested in references [8, 9]. The authors of these papers argued that the specific heat $c(T)$ of the frustrated-lattice Heisenberg model such as kagome, hyperkagome, etc., illustrates a hidden energy scale (which is considerably lower than $J$) as an extra low-temperature peak. Moreover, the accumulated entropy equals the residual ground-state entropy of the Ising model on the same lattice.
@@
In this paper, we investigate the $S=1/2$ Heisenberg model on four three-dimensional lattices, see figure 1. Two of them, simple-cubic and diamond, are bipartite ones, whereas other two, pyrochlore and hyperkagome, are geometrically frustrated ones. All these quantum spin-lattice systems are widely used in the theory of magnetism. Our goal is to discuss thermodynamic quantities for these models focusing on the specific heat in order to highlight the effect of lattice geometry in the features of $c(T)$ for the quantum Heisenberg ferro- and antiferromagnets. To this end, we use quantum Monte Carlo simulations (ferromagnets and bipartite-lattice antiferromagnets) and high-temperature expansion series analysis (frustrated-lattice antiferromagnets). We have to notice that some of the presented results are not completely new but are scattered over literature, and here we have collected them in one place accompanied by still not reported results and all of these are considered from a common perspective.
@@
The rest of the paper is organized as follows. In section 2, we introduce the models and methods. Then, in section 3, we report our findings focusing on the temperature profiles of the specific heat. Finally, we discuss and summarize in section 4. Some additional information on the entropy-method calculations for the pyrochlore and hyperkagome cases is presented in appendix A and B, respectively.
@@
Figure 1: (Colour online) Simple-cubic (upper left), diamond (upper right), pyrochlore (lower left), and hyperkagome (lower right) lattices which are considered in this paper. Different colors of discs denote different sublattices, i.e., two sublattices for the diamond lattice and four sublattices for the pyrochlore lattice. The unit cell contains 1 (simple cubic), 2 (diamond), 4 (pyrochlore), or 12 (hyperkagome) sites (magnetic atoms). For a detailed description see the main text, section 2.1.
2 Setting up the problem
2.1 Lattices
@@
Let us describe the lattices to be considered, see figure 1. Such descriptions are required for quantum Monte Carlo simulations using the ALPS package [10, 11] as well as for generating high-temperature expansion series [12].
@@
The sites on the simple cubic lattice are defined by ${\bf{R}}_{\bf{n}}=n_{x}{\bf{e}}_{x}+n_{y}{\bf{e}}_{y}+n_{z}{\bf{e}}_{z}$, where ${\bf{e}}_{x}=(1,0,0)$, ${\bf{e}}_{y}=(0,1,0)$, ${\bf{e}}_{z}=(0,0,1)$ and $n_{x}$, $n_{y}$, $n_{z}$ are integers. Here and below, we set the length of the (unit-cell) cube edge to unity.
@@
The diamond lattice consists of two interpenetrating face-centered cubic sublattices displaced by a vector $(1/4,1/4,1/4)$. That is, the sites on the diamond lattice are defined by ${\bf{R}}_{{\bf{m}}\alpha}={\bf{R}}_{\bf{m}}+{\bf{r}}_{\alpha}$. The sites of the face-centered-cubic lattice are determined by ${\bf{R}}_{\bf{m}}=m_{1}{\bf e}_{1}+m_{2}{\bf e}_{2}+m_{3}{\bf e}_{3}$, where ${\bf e}_{1}=(0,1/2,1/2)$, ${\bf e}_{2}=(1/2,0,1/2)$, ${\bf e}_{3}=(1/2,1/2,0)$ and $m_{1}$, $m_{2}$, $m_{3}$ are integers; the origins of the two face-centered-cubic sublattices are taken to be ${\bf{r}}_{1}=(0,0,0)$ and ${\bf{r}}_{2}=(1/4,1/4,1/4)$.
@@
The pyrochlore lattice consists of four interpenetrating face-centered-cubic sublattices with the lattice sites defined by ${\bf{R}}_{{\bf{m}}\alpha}={\bf{R}}_{\bf{m}}+{\bf{r}}_{\alpha}$. Here, the sites of the face-centered-cubic lattice are defined above, whereas the origins of the four face-centered-cubic sublattices are taken to be ${\bf{r}}_{1}=(0,0,0)$, ${\bf{r}}_{2}=(0,1/4,1/4)$, ${\bf{r}}_{3}=(1/4,0,1/4)$, and ${\bf{r}}_{4}=(1/4,1/4,0)$.
@@
Finally, the hyperkagome lattice has 12 sites in a cubic unit cell. The sites on the hyperkagome lattice are defined by ${\bf{R}}_{\bf{n}\alpha}={\bf{R}}_{\bf{n}}+{\bf{r}}_{\alpha}$. Here, ${\bf{R}}_{\bf{n}}=n_{x}{\bf{e}}_{x}+n_{y}{\bf{e}}_{y}+n_{z}{\bf{e}}_{z}$ generates a simple-cubic-lattice translation (see above) and the origins of twelve equivalent sublattices are defined by ${\bf{r}}_{\alpha}$, $\alpha=1,2,\dots,12$, where ${\bf{r}}_{1}=(-1/2,0,1/2)$, ${\bf{r}}_{2}=(-1/4,3/4,1/2)$, ${\bf{r}}_{3}=(-1/2,3/4,1/4)$, ${\bf{r}}_{4}=(-1/4,1/4,0)$, ${\bf{r}}_{5}=(-1/2,1/4,3/4)$, ${\bf{r}}_{6}=(-1/4,1/2,3/4)$, ${\bf{r}}_{7}=(-3/4,1/2,1/4)$, ${\bf{r}}_{8}=(0,1/2,1/2)$, ${\bf{r}}_{9}=(0,1/4,1/4)$, ${\bf{r}}_{10}=(-3/4,3/4,0)$, ${\bf{r}}_{11}=(0,0,0)$, and ${\bf{r}}_{12}=(-3/4,0,3/4)$.
@@
2.2 Methods
2.2.1 Quantum Monte Carlo
@@
2.2.2 High-temperature expansion series augmented by the entropy method
Using reference [12], we have generated the high-temperature expansion series for the specific heat (per site) $c(\beta)$, $\beta=1/T$,
$c(\beta)=\sum_{i=2}^{n}d_{j}\beta^{i}.$ (2.1)
Here, for the pyrochlore lattice $n=17$, the coefficients $d_{2},\ldots,d_{13}$ were reported in reference [15] and $d_{14},\ldots,d_{17}$ in reference [16]; for the hyperkagome lattice $n=20$ and the coefficients $d_{17},\ldots,d_{20}$,
@@ 2.2.2 High-temperature expansion series augmented by the entropy method
complement those that were reported in reference [17], see also appendix B.
@@
Within the entropy method [18, 19, 20], one works with the entropy as a function of the energy $s(e)$, $e_{0}\leqslant e\leqslant 0$, $e_{0}$ is the ground-state energy, and the Maclaurin series for $s(e)$ is known from equation (2.1). Assuming gapples low-energy excitations (they yield a power-law decay of the specific heat $c(T)$ as $T\to 0$ with the exponent $\alpha$), one interpolates an auxiliary function $G(e)=[s(e)]^{(1+\alpha)/\alpha}/(e-e_{0})$ between $e=e_{0}$ and $e=0$ using Padé approximants $[u,d](e)$ with $d+u\leqslant n$, $n=17$ (pyrochlore) or $n=20$ (hyperkagome): $G(e)\approx[(\ln 2)^{\alpha/(1+\alpha)}/(-e_{0})]\,[u,d](e)$. As a result, one gets the approximate entropy $s(e)\approx[(e-e_{0})G(e)]^{\alpha/(1+\alpha)}$ and therefore, the specific heat $c(T)$ in the parametric form: $T=1/s^{\prime}(e)$, $c(e)=-[s^{\prime}(e)]^{2}/s^{\prime\prime}(e)$; here $(\ldots)^{\prime}={\rm d}(\ldots)/{\rm d}e$. Thus, besides high-temperature expansion series (2.1), one needs the values of $e_{0}$ and $\alpha$. Under another assumption, when low-energy excitations are gapped, one works with another auxiliary function $G(e)=(e-e_{0})[s(e)/(e-e_{0})]^{\prime}\approx(\ln 2/e_{0})[u,d](e)$, which yields the approximate entropy $s(e)/(e-e_{0})\approx-\ln 2/e_{0}-\int_{e}^{0}{\rm d}\xi G(\xi)/(\xi-e_{0})$, and one needs as an input the value of $e_{0}$ only.
@@
It should be noted that even if the required input information is unavailable, one may still proceed with the entropy method and find those parameters that are missing searching for such their values which yield the largest number of close to each other $c(T)$ based on various Padé approximants $[u,d](e)$ [21]. Such a heuristic criteria were used in 2020 [15] in the study of the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet on the basis of 13th order high-temperature expansion series to arrive at $e_{0}\approx-0.52$ and $\alpha=2$ as a most favorable set of the unknown parameters. For the hyperkagome-lattice case, similar arguments lead to $e_{0}$ which lies between 0.441-0.441 and 0.435-0.435 and $\alpha=2$ [22].
@@
Since 2021, there have been several proposals for the ground-state energy of the pyrochlore-lattice $S=1/2$ Heisenberg antiferromagnet: 0.490(6)-0.490(6) (density-matrix renormalization group method [23]), 0.477(3)-0.477(3) (variational Monte Carlo methods [24]). Besides, more terms in high-temperature expansion series have become available now [16]. Furthermore, for the hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnet, four more coefficients in equation (2.1) are available now in comparison to reference [17], see equation (2.2.2). All this motivates us to reconsider the entropy-method calculations for the pyrochlore- and hyperkagome-lattice $S=1/2$ Heisenberg antiferromagnets and refine the temperature profile $c(T)$, see section 3.2.2 below.
3 Specific heat $c(T)$
3.1 Ferromagnets
Our findings for ferromagnets are collected in the middle column of table 1 and in figure 2. As is generally known, spin waves or magnons are low-lying excitations and therefore $c(T)\propto T^{3/2}$ for all ferromagnets. An order-disorder phase transition occurs at the Curie temperature $T_{\rm C}$, see table 1. Note that $T_{\rm C}$ strongly depends on the lattice geometry. Around $T_{\rm C}$, one expects the critical behavior, $c(T)-c(T_{\rm C})\propto|T-T_{\rm C}|^{-\alpha}$, $\alpha\approx-0.12$ (3D Heisenberg universality class, see reference [25]). Here, the critical exponent $\alpha<0$ and the specific heat is finite at $T_{\rm C}$. In paramagnetic phase, as the temperature tends to infinity, all quantum Heisenberg feromagnets become the Curie paramagnet. Note a peculiar behavior for the hyperkagome-lattice case: $c(T)$ exhibits a sloped shoulder in the paramagnetic phase (green curves in figure 2). The existence of the sloped shoulder in $c(T)$ is also confirmed by simple Padé approximants to the high-temperature series for $c(\beta)$ (2.1), compare $[10,10](T)$, $[9,10](T)$, and $[9,9](T)$ [indistinguishable from $[10,10](T)$] to quantum Monte Carlo data in figure 3. Meanwhile we do not have a rigorous and convincing explanation for that and this issue needs further study.